2 bananas are to be selected from a group of 8. In how many ways can this be done?
Answer by jim_thompson5910(35256)

We have two slots to fill: Slot A, Slot B

There are 8 choices for slot A

There are 7 choices for slot B

There are 8*7 = 56 permutations possible. This would be the answer if order mattered.

Since order does not matter, we must divide that result by 2 to get 56/2 = 28

We divide by two to correct for double-counting.

The final answer is 28

Side Note: we can use the combination formula to get

n C r = (n!)/(r!*(n-r)!)
8 C 2 = (8!)/(2!*(8-2)!) ... plug in n = 8 and r = 2
8 C 2 = (8!)/(2!*6!)
8 C 2 = (8*7*6!)/(2!*6!)
8 C 2 = (8*7)/(2!)
8 C 2 = (8*7)/(2*1)
8 C 2 = 56/2
8 C 2 = 28

which is an alternative way to get the same answer